# Anisotropic linear waves and breakdown of the momentum expansion in spin magnetohydrodynamics

### v1

@@ IV.1 Leading-order solutions
×$\times(\omega-\tilde{v}_{3}^{\prime})(\omega-\tilde{v}_{4}^{\prime})(\omega+i\tilde{\Gamma}).$
@@
Here, we get four propagating modes, which are known as the magneto-sonic waves. When the sign in Eq. (48a) takes plus, they are called the fast mode; In the other case, they are called the slow modes. We find that the leading-order solutions are exactly the same as the case without spin (see, e.g., Refs. [83, 87, 82] and references therein). It can be shown that $0\leq v_{1,2}\leq 1$ assuming that $0\leq c_{s}\leq 1$ and $0\leq v_{A}\leq 1$ (see Ref. [82]).
@@ IV.2 Next-to-leading order solutions
×$\times(\omega+i\Gamma_{\parallel}+i\tilde{\omega}_{3}k^{2})(\omega+i\Gamma_{\perp}+i\tilde{\omega}_{4}k^{2}),\quad$
@@
$\omega$ $=$ ±$\pm v_{A}k\cos\theta-i\tilde{\omega}_{1,2}k^{2},$ (52a)
@@
×$\times(\omega-v_{2}k+i\tilde{\omega}_{3}^{\prime}k^{2})(\omega+v_{2}k+i\tilde{\omega}_{4}^{\prime}k^{2})$
×$\times(\omega+i\Gamma_{\parallel}+i\tilde{\omega}_{5}k^{2}).$
@@
$\omega$ $=$ ±$\pm v_{1}k+ik^{2}\frac{W_{1}-W_{2}v_{1}^{2}}{2(v_{1}^{2}-v_{2}^{2})},$ (54a)
$\omega$ $=$ ±$\pm v_{2}k-ik^{2}\frac{W_{1}-W_{2}v_{2}^{2}}{2(v_{1}^{2}-v_{2}^{2})},$ (54b)
@@
One can compare the above results with the first-order solutions without spin [82]. First, the propagating modes in the $k$ expansion, which are shown in Eqs. (52a), (54a), and (54b), have the same forms as the damping Alfven and magneto-sonic waves up to the replacement of the shear viscosity $\eta_{\parallel}$ by $\tilde{\eta}_{\parallel}$ defined in Eq. (51). In other words, only when there is a nonzero cross rotational viscosity $\xi_{\parallel}$, the propagating modes are modified by dynamics of angular momentum. In Eq. (16), the cross rotational viscosity converts the antisymmetric perturbation to the symmetric stress, and modifies the Alfven and sound modes.
@@ IV.3 Breakdown of the small-momentum expansion
We have obtained the solutions for an arbitrary angle $\theta$ up to the $k^{2}$ order in Sec. IV.2. Here, we, however, point out that the small-$k$ expansion breaks down depending on the angle $\theta$, and thus provide an alternative method of expansion in such an angle regime. One can realize the breakdown of the small-$k$ expansion by comparing the solutions in Eqs. (52) and (54) with the solutions at specific angles shown in Appendix. A; The latter solutions can be obtained without any expansion.
@@
$\omega$ $=$ ±$\pm\sqrt{v_{A}^{2}k^{2}\cos^{2}\theta-\frac{1}{4}(\tilde{\rho}-\tilde{\eta})^{2}k^{4}}-\frac{i}{2}(\tilde{\rho}+\tilde{\eta})k^{2},$ (58)
where $\tilde{\rho}=\rho^{\prime}_{\perp}\cos^{2}\theta+\rho^{\prime}_{\parallel}\sin^{2}\theta$ and $\tilde{\eta}=\frac{1}{h}(\eta_{\parallel}\cos^{2}\theta+\eta_{\perp}\sin^{2}\theta)$. This is the solution without any expansion. The square root in Eq. (58) becomes a pure imaginary number when $k$ grows larger than $2v_{A}\left|\frac{\cos\theta}{\tilde{\rho}-\tilde{\eta}}\right|$. Namely, the dispersion relation loses the real part, turning to a purely diffusive mode. Clearly, this behavior is different from that of the solution in the small-$k$ expansion that always has the Alfven velocity as the group velocity. In this regim where $k>2v_{A}\left|\frac{\cos\theta}{\tilde{\rho}-\tilde{\eta}}\right|$, one can instead expand Eq. (58) with $\cos\theta$ near $\cos\theta=0$. This alternative expansion is valid in the regime where the square root is imaginary.
@@
where $v_{f}^{2}=c_{s}^{2}+v_{A}^{2}(1-c_{s}^{2})$. However, we notice that the limiting behavior of the two slow magneto-sonic modes (60b) disagree with Eq. (67) and the second solution in Eq. (68a) that are obtained without any expansion. This disagreement also stems from the existence of a non-analytic boundary in the $k$-$\theta$ plane as in the case of the Alfven wave. The correct result at this limit is obtained in Eq. (A.2) by taking $\theta=\pi/2$ without the $k$ expansion.
@@
These solutions in Eqs. (59) and (61) are expected to be all stable as seen in Figs. 1 and 2. According to the solutions in Appendix A.2, one can easily confirm that all the modes are damping at $\theta=\pi/2$. As we decrease $\theta$, the imaginary parts of the solutions approach those of the Alfven and magneto-sonic modes that have the definite signs as shown in Sec. IV.2. Therefore, the solutions in Eqs. (59) and (61) are stable (unless there arise non-monotonic behaviors with respect to $\theta$).
@@
We stress that the above analytic solutions are obtained for an arbitrary angle $\theta$ in the small-momentum expansion. However, these solutions indicate that the small-momentum expansion is spoiled due to the blow up of the higher-order terms when the angle $\theta$ approaches the right angle where the linear waves propagate in nearly perpendicular to an equilibrium magnetic field. We showed that there is a critical angle where the two competing expansion parameters, the momentum $k$ and $\cos\theta$, have similar magnitudes. Since the critical angle depend on the momentum, the solution has a non-analytic boundary in the $k-\theta$ plane. On the other side of the boundary, we provided the alternative solution in the small-cosine expansion. We confirmed that the momentum and cosine expansions well reproduce the numerical solutions in the appropriate domains separated by the boundary. We point out that this non-analytic structure is a prevalent issue in anisotropic systems.
We would like to give outlook. To address the splitting between $\Lambda$ and $\bar{\Lambda}$ spin polarization in the framework of spin MHD, it is important to include a finite chemical potential since a net spin polarization is not induced in a charge neutral fluid. Then, one may need to include a finite electric current and an electric field accordingly. They are quasihydrodynamic modes, and require an extended framework. Also, it is interesting to investigate the spin relaxation to a nonzero thermal vorticity. One needs to include a zeroth-order vorticity into the constitutive equations (cf. Ref. [29]) and perform the linear-mode analysis near a vortical equilibrium configuration. Last but not least, extensions to the Israel-Stewart type frameworks will be useful for numerical implementation toward applications to relativistic-heavy ion collisions.
@@ Acknowledgements.
The authors thank Yihui Tu, Xu-Guang Huang and Yi Yin for useful comments. This work is partially supported by the JSPS KAKENHI under grant Nos. 20K03948 and 22H01216.
@@ A.1 Eigenmodes at $k_{\perp}=0$
00 0 $=$ $\left(\begin{array}[]{ccc|ccc}Bk_{\parallel}&\omega+i\rho_{\perp}^{\prime}k_{\parallel}^{2}&0&&\\ h\omega+i(\eta_{\parallel}-2\xi_{\parallel}+\gamma_{\parallel})k_{\parallel}^{2}&\frac{h}{B}v_{A}^{2}k_{\parallel}&\frac{4}{\chi}(\gamma_{\parallel}-\xi_{\parallel})k_{\parallel}&&\\ -2(\gamma_{\parallel}-\xi_{\parallel})k_{\parallel}&0&\omega+i\Gamma_{\parallel}&&\\ \hline\cr&&&\omega&&-h(1-v_{A}^{2})k_{\parallel}\\ &&&-c_{s}^{2}k_{\parallel}&&h(1-v_{A}^{2})\omega+i\zeta_{\parallel}k_{\parallel}^{2}\end{array}\right)\begin{pmatrix}\delta u_{x}\\ \delta B_{x}\\ \delta\sigma_{y}\\ \delta\epsilon\\ \delta u_{z}\end{pmatrix},$ (62f)
00 0 $=$ $\left(\begin{array}[]{ccc|c}Bk_{\parallel}&\omega+i\rho_{\perp}^{\prime}k_{\parallel}^{2}&0&\\ h\omega+i(\eta_{\parallel}-2\xi_{\parallel}+\gamma_{\parallel})k_{\parallel}^{2}&\frac{h}{B}v_{A}^{2}k_{\parallel}&-\frac{4}{\chi}(\gamma_{\parallel}-\xi_{\parallel})k_{\parallel}&\\ 2(\gamma_{\parallel}-\xi_{\parallel})k_{\parallel}&0&\omega+i\Gamma_{\parallel}&\\ \hline\cr&&&\omega+i\Gamma_{\perp}\end{array}\right)\begin{pmatrix}\delta u_{y}\\ \delta B_{y}\\ \delta\sigma_{x}\\ \delta\sigma_{z}\end{pmatrix}\,.$ (62k)
@@
±$\pm v_{A}k_{\parallel}-\frac{i}{2}\Big{(}\rho^{\prime}_{\perp}+\frac{\eta_{\parallel}\gamma_{\parallel}-\xi_{\parallel}^{2}}{h\gamma_{\parallel}}\Big{)}k_{\parallel}^{2}.$
@@
$\omega$ $=$ ±$\pm c_{s}k_{\parallel}-\frac{i}{2}\zeta^{\prime}_{\parallel}k_{\parallel}^{2},$ (64b)
where $\zeta^{\prime}_{\parallel}=\frac{\zeta_{\parallel}}{h(1-v_{A}^{2})}$. Again, the spin mode is damped out in the hydrodynamic limit with $\Gamma_{\perp}$ from the non-vanishing right-hand side in Eq. (3). The other two modes are sound waves that are damped out due to one of the bulk viscosities. All those modes have longitudinal polarizations with respect to the momentum.

### v2

@@ IV.1 Leading-order solutions
[…]
@@
Here, we get four propagating modes, which are known as the magneto-sonic waves. When the sign in Eq. (50a) takes plus, they are called the fast modes; In the other case, they are called the slow modes. We find that the leading-order solutions are exactly the same as the case without spin (see, e.g., Refs. [91, 99, 90] and references therein). It can be shown that $0\leq v_{1,2}\leq 1$ assuming that $0\leq c_{s}\leq 1$ and $0\leq v_{A}\leq 1$ (see Ref. [90]).
@@ IV.2 Next-to-leading order solutions
$\times(\omega+i\Gamma_{\parallel}+i\tilde{\omega}_{3}k^{2})(\omega+i\Gamma_{\perp}+i\tilde{\omega}_{4}k^{2}),\quad$
@@
$\omega$ $=$ $\pm v_{A}k\cos\theta-i\tilde{\omega}_{1,2}k^{2},$ (54a)
@@
$\times(\omega-v_{2}k+i\tilde{\omega}_{3}^{\prime}k^{2})(\omega+v_{2}k+i\tilde{\omega}_{4}^{\prime}k^{2})$
$\times(\omega+i\Gamma_{\parallel}+i\tilde{\omega}_{5}k^{2}).$
@@
$\omega$ $=$ $\pm v_{1}k+ik^{2}\frac{W_{1}-W_{2}v_{1}^{2}}{2(v_{1}^{2}-v_{2}^{2})},$ (56a)
$\omega$ $=$ $\pm v_{2}k-ik^{2}\frac{W_{1}-W_{2}v_{2}^{2}}{2(v_{1}^{2}-v_{2}^{2})},$ (56b)
@@
One can compare the above results with the first-order solutions without spin [90]. First, the propagating modes in the $k$ expansion, which are shown in Eqs. (54a), (56a), and (56b), have the same forms as the damping Alfven and magneto-sonic waves up to the replacement of the shear viscosity $\eta_{\parallel}$ by $\tilde{\eta}_{\parallel}$ defined in Eq. (53). In other words, only when there is a nonzero cross rotational viscosity $\xi_{\parallel}$, the propagating modes are modified by dynamics of angular momentum. In Eq. (18), the cross rotational viscosity converts the antisymmetric perturbation to the symmetric stress, and modifies the Alfven and sound modes. The other three modes in Eqs. (54b), (54c), and (56c) are damping spin modes. The damping factor remains in the zero momentum limit, and spin density does not persist in the strict equilibrium limit because spin angular momentum is not a conserved quantity by itself [22].
@@ IV.3 Breakdown of the small-momentum expansion
We have obtained the solutions for an arbitrary angle $\theta$ up to the $k^{2}$ order in Sec. IV.2. Here, however, we point out that the small-$k$ expansion breaks down depending on the angle $\theta$, and thus provide an alternative method of expansion in such an angle regime. One can realize the breakdown of the small-$k$ expansion by comparing the solutions in Eqs. (54) and (56) with the solutions at specific angles shown in Appendix. A. The latter solutions can be obtained without any expansion.
When $\theta=\pi/2$, the solution (54a) does not agree with any of the solutions in the limit $\theta=\pi/2$. [^2] Consistent cases can be confirmed as follows. For the set of solutions (54) containing the Alfven modes, the solutions (54) are consistent with those in Eq. (69) and Eq. (70) when $\theta=0$. When $\theta=\pi/2$, the solutions (54b) and (54c) are still consistent with the one in Eq. (74d) and the first one in Eq. (74b). Therefore, one should investigate the behavior of the solution near $\theta=\pi/2$ more closely. As discussed below, the exact solution (without the small-$k$ expansion) has a non-analytic boundary at $k=k_{c}$, where $k_{c}$ depends on $\theta$. The $k$ expansion gets invalid when $k$ goes beyond this boundary. The critical momentum $k_{c}$ approaches zero when $\cos\theta$ approaches zero. Accordingly, the Taylor expansion for $k$ breaks down when $\cos\theta$ approaches zero. This non-analytic property arises from that of a square root. Since waves propagating in opposite directions appear in pairs, the secular equation is an equation for $\omega^{2}$, and its solution for $\omega$ is in general given by a square root.
$\omega$ $=$ $\pm\sqrt{v_{A}^{2}k^{2}\cos^{2}\theta-\frac{1}{4}(\tilde{\rho}-\tilde{\eta})^{2}k^{4}}-\frac{i}{2}(\tilde{\rho}+\tilde{\eta})k^{2},$ (60)
@@
where $v_{f}^{2}=c_{s}^{2}+v_{A}^{2}(1-c_{s}^{2})$. However, we notice that the limiting behavior of the two slow magneto-sonic modes (62b) disagrees with Eq. (73) and the second solution in Eq. (74a) that are obtained without any expansion. This disagreement also stems from the existence of a non-analytic boundary in the $k$-$\theta$ plane as in the case of the Alfven wave. The correct result at this limit is obtained in Eq. (A.2) by taking $\theta=\pi/2$ without the $k$ expansion.
@@
These solutions in Eqs. (61) and (63) are expected to be stable as seen in Figs. 1 and 2. According to the solutions in Appendix A.2, one can easily confirm that all the modes are damping at $\theta=\pi/2$. As we decrease $\theta$, the imaginary parts of the solutions approach those of the Alfven and magneto-sonic modes that have the definite signs as shown in Sec. IV.2. Therefore, the solutions in Eqs. (61) and (63) are stable (unless there arise non-monotonic behaviors with respect to $\theta$).
IV.4 Extension to the Israel-Stewart equations
In the first-order hydrodynamics, diffusive modes typically lead to violation of causality as diffusion equations imply instantaneous signal propagation. Since a prescription for this issue was provided by Israel and Stewart [106, 107, 108], causality has been one of the important topics in relativistic hydrodynamics and is still a hot topic for discussion (see, e.g., Refs. [109, 110, 111, 112, 113, 114, 115] for recent developments).
Our first-order solutions, shown in Figs. 1 and 2, contain non-propagating regimes near the vertical limit, i.e., $\theta\rightarrow\pi/2$. We investigate these diffusive modes based on the IS equations by introducing a finite relaxation time $\tau$. We will find that, near the vertical limit, there remain diffusive modes in the IS equations in a low-momentum region, while they are modified to become propagating modes in the higher momentum region. Therefore, the IS-modified dispersion relation exhibits a diffusive window in the intermediate momentum region as shown below.
We focus on the Alfven modes to demonstrate how the diffusive modes, emerging near the vertical limit, are modified in the IS equations. For simplicity, we assume that the transport coefficients take isotropic values, i.e., $\eta:=\eta_{\parallel}=\eta_{\perp}$ and $\rho:=\rho_{\parallel}=\rho_{\perp}$, since this simplification does not affect a conclusion regarding the presence of residual diffusive modes in the IS equations at the qualitative level; We also suppress the spin sector for the same reason. Then, the IS equations for the Alfven modes read
$\partial_{0}\delta B_{y}=B\partial_{z}\delta u_{y}+\partial_{x}\delta\tilde{\mathcal{F}}_{xy}-\partial_{z}\delta\tilde{\mathcal{F}}_{yz},$ (64a)
$\tau\partial_{0}\delta\tilde{\mathcal{F}}_{xy}+\delta\tilde{\mathcal{F}}_{xy}=\frac{1}{\mu_{m}}{\rho}\partial_{x}\delta B_{y},$ (64b)
$\tau\partial_{0}\delta\tilde{\mathcal{F}}_{yz}+\delta\tilde{\mathcal{F}}_{yz}=-\frac{1}{\mu_{m}}{\rho}\partial_{z}\delta B_{y},$ (64c)
$h{\partial_{0}}{\delta u^{y}}=B\partial_{z}\delta B_{y}+\partial_{x}\delta\pi_{xy}+\partial_{z}\delta\pi_{yz},$ (64d)
$\tau\partial_{0}\delta\pi_{xy}+\delta\pi_{xy}={\eta}\partial_{x}\delta u_{y},$ (64e)
$\tau\partial_{0}\delta\pi_{yz}+\delta\pi_{yz}={\eta}\partial_{z}\delta u_{y}.$ (64f)
where we denote the dynamical variables in the IS equations as $\pi_{ij}$ and $\tilde{\mathcal{F}}_{ij}$ and the corresponding relaxation times as $\tau_{R}$ and $\tau_{B}$, respectively. We take both relaxation times to be the same value $\tau$. When $\tau\rightarrow 0$, these equations go back the first-order equations (39) discussed in the previous sections (without spin).
We first provide analytic solutions at the vertical limit, which are obtained as
$\omega=-\frac{ih\pm\sqrt{4\eta hk^{2}\tau-h^{2}}}{2h\tau}\,,\quad-\frac{i\mu_{m}\pm\sqrt{4\mu_{m}\rho k^{2}\tau-\mu_{m}^{2}}}{2\mu_{m}\tau}\,.$ (65)
The other two equations for $\delta\tilde{\mathcal{F}}_{yz}$ and $\delta\pi_{yz}$ are decoupled in this limit. Expanding these solutions for a small $k$, one finds two diffusive modes and two gapped modes
$\omega$ $=$ $-i\frac{\eta}{h}k^{2}\,,\quad-\frac{i}{\tau}+i\frac{\eta}{h}k^{2}\,$, (66a)
$\omega$ $=$ $-i\frac{\rho}{\mu_{m}}k^{2}\,,\quad-\frac{i}{\tau}+i\frac{\rho}{\mu_{m}}k^{2}\,.$ (66b)
Two of these modes, which depend on $\tau$, are gapped modes introduced by the IS relaxation equations. The other two gapless diffusive modes are independent of the relaxation time at the leading order in $k$ and reduce to the Alfven modes in the first-order solutions in Fig. 1 when $\tau\to 0$; Thus, these modes are identified as the “IS-modified Alfven modes.” In Panel 1 of Fig. 3, blue curves show the IS-modified Alfven modes, while orange curves show the the Alfven modes in the first-order solutions at $\tau=0$. We take $\eta h^{-3/4}=0.2$, $\rho h^{1/4}=0.2$, $\tau h^{1/4}=1$, and $v_{A}=0.7$, where dimensionful quantities are rescaled by $h$ in Eq. (30).
The small-$k$ expansion (66b) shows that the IS-modified Alfven modes remain diffusive modes at the vertical limit even after the IS relaxation equations are introduced. This behavior is in clear contrast to the typical behavior of dispersion relations from the IS equations where diffusive modes in the first-order equations are modified to become propagating modes in a small momentum region down to vanishing $k$ (see, e.g., Ref. [116]). As we increase $k$, the square-root arguments in the full solution (65) turn into positive values at the critical momenta
$k_{c}^{\rm IS}=\sqrt{\frac{h}{4\eta\tau}}\,,\quad\sqrt{\frac{\mu_{m}}{4\rho\tau}}\,.$ (67)
At each of these critical momenta, a diffusive mode merges with one of the gapped relaxation modes to form a pair of propagating modes as shown with the blue curves in the same figure (Panel 1 of Fig. 3).
Figure 3: Dispersion relations of the IS-modified Alfven modes (blue curves) and the first-order Alfven modes (orange curves) at fixed angles. Solid and dotted curves show the real and imaginary parts of the dispersion relations. See also Fig. 4 for general angles.
Figure 4: Unshaded and shaded regions show diffusive and propagating regimes, respectively, in the IS-modified Alfven modes (blue shade) and the first-order Alfven modes (orange shade). The IS-modified Alfven modes retain a diffusive window in the unshaded “shark-fin” region.
On top of the above observations at the vertical limit, we examine the angle dependence of the dispersion relations. In Fig. 4, diffusive and propagating regimes are shown by the unshaded and shaded regions, respectively, as a phase diagram with respect to $\theta$ and $k$. The numbered cross-sections at different angles correspond to the dispersion relations shown in Fig. 3. We take the same parameters in Figs. 3 and 4.
First, let us examine the low-momentum regions. In case of the first-order MHD discussed in the previous sections, we have the propagating Alfven waves as we decrease the angle from the vertical limit, which corresponds to the region shaded in dark orange. However, the small-momentum expansion breaks down at the critical momentum $k_{c}(\theta)$ discussed in Sec. IV.3; The Alfven modes in the first-order MHD become diffusive above $k_{c}(\theta)$. As shown in Fig. 4, the IS-modified Alfven modes basically succeed these properties, though the critical momentum is slightly modified as we find a slight deviation between the orange and blue boundaries. These propagating modes are also confirmed in the low-momentum regions in Panels 2, 3, and 4 of Fig. 3.
In the high-momentum region, diffusive modes are modified by the relaxation dynamics introduced in the IS equations in general. Indeed, as already discussed above at the vertical limit, the diffusive regime in the first-order solutions turns into the propagating regime. This modification is found in all the angles in Fig. 3 in the high-momentum regions. Recall, however, that this modification is effective only above the other critical momentum $k_{c}^{\rm IS}$ defined in Eq. (67) that is controlled by the relaxation time $\tau$ and depends on $\theta$ in general.
It is, therefore, interesting to focus on a momentum window defined by the two critical momenta, $k_{c}(\theta)\leq k\leq k_{c}^{\rm IS}(\theta)$. Since the critical momenta depend on the angle $\theta$, the window is closed when $k_{c}(\theta)=k_{c}^{\rm IS}(\theta)$. At this angle, the low- and high-momentum propagating regimes touch each other (see Panel 3 in Fig. 3). Above this angle (closer to the vertical limit), a diffusive window persists even after a finite relaxation time is introduced with the IS equations. This regime corresponds to the unshaded “shark-fin” region in Fig. 4. The presence of the diffusive window in an intermediate momentum region deserves further study in future works.
@@ IV.4 Extension to the Israel-Stewart equations
We stress that the above analytic solutions are obtained for an arbitrary angle $\theta$ in the small-momentum expansion. However, these solutions indicate that the small-momentum expansion is spoiled due to the blow up of the higher-order terms when the angle $\theta$ approaches the right angle where the linear waves propagate nearly perpendicular to an equilibrium magnetic field. We showed that there is a critical angle where the two competing expansion parameters, the momentum $k$ and $\cos\theta$, have similar magnitudes. Since the critical angle depends on the momentum, the solution has a non-analytic boundary in the $k$-$\theta$ plane. This boundary defines the critical momentum $k_{c}(\theta)$ when we consider the dispersion relations as functions of $k$ at a fixed $\theta$. On the other side of the boundary, we provided the alternative solution in the small-cosine expansion. We confirmed that the momentum and cosine expansions well reproduce the numerical solutions in the appropriate domains separated by the boundary. We point out that this non-analytic structure is a prevalent issue in anisotropic systems.
We also included the relaxation dynamics for the electric resistivity and shear viscosity with the IS framework. These higher-order corrections modify the dispersion relations in a high momentum region. In particular, the purely diffusive regime, which was found in the first-order solutions above the critical momentum $k_{c}(\theta)$, turns into a propagating regime due to the relaxation dynamics. The boundary between the diffusive and propagating regimes defines another critical momentum $k_{c}^{\rm IS}(\theta)$. We found that there remains a diffusive window between the lower critical momentum $k_{c}$ and the upper critical momentum $k_{c}^{\rm IS}$ as shown in Fig. 4.
We would like to give a future outlook. To address the splitting between $\Lambda$ and $\bar{\Lambda}$ spin polarization in the framework of spin MHD, it is important to include a finite chemical potential since a net spin polarization is not induced in a charge neutral fluid. Then, one may need to include a finite electric current and an electric field accordingly. They are quasihydrodynamic modes, and require an extended framework. Also, it is interesting to investigate the spin relaxation to a nonzero thermal vorticity. One needs to include a zeroth-order vorticity into the constitutive equations (cf. Ref. [29]) and perform the linear-mode analysis near a vortical equilibrium configuration. Last but not least, extensions to the Israel-Stewart type frameworks will be useful for numerical implementation toward applications to relativistic-heavy ion collisions.
@@ Acknowledgements.
The authors thank Yi-hui Tu, Xu-Guang Huang, and Yi Yin for useful comments. This work is supported by the NSFC under grant No.12505149, and the JSPS KAKENHI under grant No.20K03948 and No.22H01216.
@@ A.1 Eigenmodes at $k_{\perp}=0$
$0$ $=$ $\left(\begin{array}[]{ccc|ccc}Bk_{\parallel}&\omega+i\rho_{\perp}^{\prime}k_{\parallel}^{2}&0&&\\ h\omega+i(\eta_{\parallel}-2\xi_{\parallel}+\gamma_{\parallel})k_{\parallel}^{2}&\frac{h}{B}v_{A}^{2}k_{\parallel}&\frac{4}{\chi}(\gamma_{\parallel}-\xi_{\parallel})k_{\parallel}&&\\ -2(\gamma_{\parallel}-\xi_{\parallel})k_{\parallel}&0&\omega+i\Gamma_{\parallel}&&\\ \hline\cr&&&\omega&&-h(1-v_{A}^{2})k_{\parallel}\\ &&&-c_{s}^{2}k_{\parallel}&&h(1-v_{A}^{2})\omega+i\zeta_{\parallel}k_{\parallel}^{2}\end{array}\right)\begin{pmatrix}\delta u_{x}\\ \delta B_{x}\\ \delta\sigma_{y}\\ \delta\epsilon\\ \delta u_{z}\end{pmatrix}$, (68f)
$0$ $=$ $\left(\begin{array}[]{ccc|c}Bk_{\parallel}&\omega+i\rho_{\perp}^{\prime}k_{\parallel}^{2}&0&\\ h\omega+i(\eta_{\parallel}-2\xi_{\parallel}+\gamma_{\parallel})k_{\parallel}^{2}&\frac{h}{B}v_{A}^{2}k_{\parallel}&-\frac{4}{\chi}(\gamma_{\parallel}-\xi_{\parallel})k_{\parallel}&\\ 2(\gamma_{\parallel}-\xi_{\parallel})k_{\parallel}&0&\omega+i\Gamma_{\parallel}&\\ \hline\cr&&&\omega+i\Gamma_{\perp}\end{array}\right)\begin{pmatrix}\delta u_{y}\\ \delta B_{y}\\ \delta\sigma_{x}\\ \delta\sigma_{z}\end{pmatrix}\,$. (68k)
@@
$\pm v_{A}k_{\parallel}-\frac{i}{2}\Big(\rho^{\prime}_{\perp}+\frac{\eta_{\parallel}\gamma_{\parallel}-\xi_{\parallel}^{2}}{h\gamma_{\parallel}}\Big)k_{\parallel}^{2}.$
@@
$\omega$ $=$ $\pm c_{s}k_{\parallel}-\frac{i}{2}\zeta^{\prime}_{\parallel}k_{\parallel}^{2},$ (70b)
where $\zeta^{\prime}_{\parallel}=\frac{\zeta_{\parallel}}{h(1-v_{A}^{2})}$. Again, the spin mode is damped out in the hydrodynamic limit with $\Gamma_{\perp}$ from the non-vanishing right-hand side in Eq. (3). The other two modes are sound waves that are damped out due to one of the bulk viscosities. All these modes have longitudinal polarizations with respect to the momentum.